OpenStudy (anonymous):

Solve B = –4(c – 2d) for d.

OpenStudy (vishweshshrimali5):

well i can't solve it

OpenStudy (vishweshshrimali5):

sorry i mean i can solve it

OpenStudy (vishweshshrimali5):

See B = –4(c – 2d) => B = -4(c-3d) => B = -4c + 12d => B + 4c = 12 d => d = (B+4c)/12 This is your answer Satisfied ?